OpenStudy (anonymous):

with a short time remaining in the day a ups driver has time to make deliveries at only 4 locations among the 10 locations remaining. How many different routes are there if the order of the deliveries must be taken into account??

OpenStudy (noahbred):

is this multiplication?

OpenStudy (anonymous):

It's stats.

OpenStudy (anonymous):

@noahbred

OpenStudy (dumbcow):

\[10 P 4 = \frac{10!}{6!} = 10*9*8*7\]

OpenStudy (anonymous):

So what's the answer from that

OpenStudy (anonymous):

5040?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yes 5040 ... i figured you can do multiplication on your own

OpenStudy (anonymous):

Yup thanks can you help me with one more ;)

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

How many different routes are there if the order of the deliveries must be taken into account and there are 8 locations remaining?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

same thing just change the 4 to an 8

OpenStudy (anonymous):

90? :)

OpenStudy (dumbcow):

here is general rule for permutations \[nPr = \frac{n!}{(n-r)!}\]

OpenStudy (anonymous):

so its not 90?

OpenStudy (dumbcow):

no not 90, that would be if 2 remaining

OpenStudy (anonymous):

wait i got 75,600???

OpenStudy (anonymous):

I am confused help me

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

1814400

OpenStudy (anonymous):

?

OpenStudy (dumbcow):

ok 8 remaining means we have 10P8 permutations --> 10*9*8*7*6*5*4*3

OpenStudy (anonymous):

was my last answer correct?

OpenStudy (dumbcow):

it looks right, i haven't done the multiplying

OpenStudy (anonymous):

Can you help me with one more thing ;)

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

a player wins first prize by selecting the correct 5-number combination when 5 different numbers from 1 through 59 are drawn. It doesn’t matter what order the numbers are drawn in as long as all 5 numbers match - How many 5-number combinations are there out of 59 numbers - How many 5-number combinations would there be if they were drawn from 60 different numbers -------------------------------- is n= 59 r= 5?

OpenStudy (anonymous):

@dumbcow